If Sally can paint a house in 4 hours and John can paint the same house in 6 hours, how long will it take to paint the same house together?
A. 2 hr. 24 min
B.3 hr. 12 min.
C. 3 hr. 44 min.
D. 4 hr. 10 min.
E. 4 hr. 33 min.
I tried to solve by dividing both of their times by 2 but that gave me 5 and it wasn't one of the answer choices. How would you solve this?

Found 2 solutions by stanbon, dolphinlover100:
Answer by stanbon(75887)

If Sally can paint a house in 4 hours and John can paint the same house in 6 hours, how long will it take to paint the same house together?
---------
Sally DATA:
Time = 4 hrs/job ; Rate = 1/4 job/hr
-----------
John DATA:
Time = 6 hrs/job ; Rate = 1/6 job/hr
----------------
Together DATA:
Time = x hrs/job ; Rate = 1/x job/hr
------------------------
EQUATION:
rate + rate = together rate
1/4+ 1/6 = 1/x
Multiply thru by 12x to get:
3x + 2x = 12
5x = 12
x = 2 hrs. (2/5)hr
x = 2 hrs 24 minutes
===============
Cheers,
Stan H.

Answer by dolphinlover100(1)

a)2 hrs and 24 min.



x x
24 ( _ + _ = 1)
4 6
6x+4x= 24
10x= 24
then just divide out